OpenStudy (anonymous):

please help Write the standard equation for the circle that passes through the points (- 5, 0), (0, 4), and (2, 4)

OpenStudy (harsimran_hs4):

assume a general eq x^2 + y^2 + 2gx + 2fy + c = 0 g , f , c are unknowns and 3 conditions/ points are given so can you find out g, f and c?

OpenStudy (anonymous):

4x^2+4y^2-79/16x+95/8-2y but its wrong

OpenStudy (anonymous):

@harsimran_hs4

OpenStudy (anonymous):

A full answer with graph can be found here: http://www.wolframalpha.com/input/?i=Equ … So how do you set it up? (x–h) ² + (y–k)² = r² we have 3 unknowns and need to solve for h, k and r. Just plug in the three points one at a time and get: (–5–h)² + (0– k)² = r² → h²+10h + 25 + k² = r² (0–h)² + (4– k)² = r² → h² +k² –8k + 16 = r² (2–h)² + (4– k)² = r² → h² – 4h + 4 + k² –8k + 16= r² All of these reduce to: h²+10h + 25 + k² = r² h² +k² –8k + 16 = r² h² – 4h + k² –8k + 20= r² If you solve the system (not a trivial task), you will get (x-1)²+(y+19/8)² = 2665/64

OpenStudy (harsimran_hs4):

so you can have a look at the above solution if still any confusion you can ask

OpenStudy (anonymous):

these three ar my unknowns -79/16x+95/8-2y

OpenStudy (anonymous):

@dusty how can i find the final answer?

OpenStudy (harsimran_hs4):

@begzat when you solved what values did you get for g, f, c?

OpenStudy (anonymous):

-79/16x + 95/8 -2y this are my solution that i get

OpenStudy (harsimran_hs4):

how come x comes in ? g = ? f = ? c = ? can you fill the blanks?

OpenStudy (anonymous):

no

OpenStudy (harsimran_hs4):

then perhaps you should first have a look at dusty`s post look at it line by line and try to figure out he has taken the general eq in another form

OpenStudy (anonymous):

hey im a girl lol :)

OpenStudy (anonymous):

im stock i have no idea

OpenStudy (harsimran_hs4):

@dusty i apologize for it

OpenStudy (anonymous):

its okay :)

OpenStudy (anonymous):

are this wrong for these points -79/16x + 95/8 -2y or not

OpenStudy (anonymous):

i mean the answer its wrong

OpenStudy (anonymous):

@dusty can you solve for me please

OpenStudy (anonymous):

why dont you do this one :) equation of circle is (x-a)^2 +(y-b)^2 = r^2 substitute A(- 5, 0) B(0, 4) C(2, 4) into circle equation and you get a system of 3 equation in a,b, r an find center and radius

OpenStudy (anonymous):

do you understand stand?

OpenStudy (anonymous):

so i have to take just one of the points

OpenStudy (anonymous):

(5 - 2)^2 + (4 - 0)^2 = r^2 3^2 + 4^2 = r^2 9 + 16 = r^2 25 = r^2 r = 5

OpenStudy (anonymous):

x^2+y^2-4x-25 = 0 is this final answer